# Optimal navigation in two-dimensional regular and turbulent flows

### v1

@@ IV.4 Example 3: Taylor-Green vortices
Next, we consider the Taylor-Green flow, see Section LABEL:sec:Ex_C. Now, the optimal control equations must be solved numerically. To define the problem over a fixed integration interval, it is convenient to introduce a rescaled time variable $\tau=t/T$, such that $\tau\in[0,1]$. Using this transformation, the governing equations take the form
$\dot{X}_{\tau}=u_{0}T\sin(kX_{\tau})\cos(kY_{\tau})+V_{0}T\cos(\theta_{\tau}),$ (21)
$\dot{Y}_{\tau}=-u_{0}T\cos(kX_{\tau})\sin(kY_{\tau})+V_{0}T\sin(\theta_{\tau}),$ (22)
$\dot{\theta}_{\tau}=u_{0}kT\big[\cos(kX_{\tau})\cos(kY_{\tau})\sin(2\theta_{\tau})$ (23)
$+\sin(kX_{\tau})\sin(kY_{\tau})\big],$
with the boundary conditions $(X_{0},Y_{0})=(x_{A},y_{A})$ and $(X_{1},Y_{1})=(x_{B},y_{B})$.
@@
The system of equations can be numerically integrated starting from the initial position $\bm{r}_{A}$ for a prescribed initial steering angle $\theta_{0}$ and total travel time $T$. However, these two parameters are unknown and must be adjusted so that the final position $(X_{1},Y_{1})$ matches the desired target point $\bm{r}_{B}$. To achieve this, we employ a shooting method followed by fine-tuning optimization. Initially, approximate values of $\theta_{0}$ and $T$ are chosen to bring the trajectory close to the target. These parameters are then refined to minimize the discrepancy
[…]
@@
between the simulated and desired endpoints. For optimization, we use gradient- and Hessian-based methods. In particular, our implementation first applies the Adam optimizer [kingma2014adam], followed by the BFGS algorithm [nocedal2006numerical]. Efficient computation of the loss gradients $\partial{\mathcal{L}}/\partial\theta_{0}$ and $\partial{\mathcal{L}}/\partial\theta_{0}$ requires the forward integration scheme to support automatic differentiation [baydin2018automatic]. Moreover, the integration must be performed with high numerical accuracy. Convergence tests with progressively reduced tolerances are essential to ensure that the solution is independent of the integration precision.
[…]
For the numerical parameters specified in Section LABEL:sec:Ex_C, we obtain:
$label{eq:opt-control-taylor-green}\theta_{0}\approx 1.56079,\quad T\approx 12.2038.$ (25)
@@
The corresponding optimal trajectory is shown in Fig. LABEL:fig:4 by a white line. The figure also illustrate the system’s sensitivity by depicting trajectories perturbed by small deviations in the initial heading $\theta_{0}\pm\pi/9000$ and by displacements $\Delta=10^{-3}$ in the initial $x_{A}$ and $y_{A}$ positions. Although these perturbations are considerably smaller than those examined in previous examples, the resulting trajectories fail to reach the target $\bm{r}_{B}$.
@@
To make a more quantitative stability analysis of the optimal trajectory, we consider the evolution of infinitesimal perturbations,
$(\delta X_{t},\delta Y_{t},\delta\theta_{t})^{T}=\hat{W}(t)(\delta X_{0},\delta Y_{0},\delta\theta_{0})^{T},$ (26)
@@
where $\hat{W}(t)$ is the linearized evolution matrix that can be obtained numerically. A polar decomposition $\hat{W}(t)=\hat{V}(t)\hat{R}(t)$ allows us to separate the effects of rotation, represented by the orthogonal matrix $\hat{R}(t)$, from those of stretching, characterized by the positive-definite matrix $\hat{V}(t)$. The eigenvalues $h_{i}(t)$ of $\hat{V}(t)$ (equal to singular values of $\hat{W}(t)$) quantify the magnitude of stretching along different directions in phase space and are used to compute the finite-time Lyapunov exponents (FTLEs), defined as
@@
The corresponding results are presented in Fig. LABEL:fig:5. The presence of a positive maximal FTLE indicates that the optimal trajectory is unstable with respect to small perturbations on the time scales needed to reach the target. This instability limits the practical applicability of the pre-computed optimal control strategy, as minor deviations in initial conditions can lead to significant trajectory divergence. Moreover, the resulting sensitivity makes it difficult to find suitable initial guess for parameters $\theta_{0}$ and $T$, that lead to a successful solution of the optimization problem and finding optimal trajectories that reach the target. Despite these challenges, computing optimal trajectories can still be practically valuable if the agent additionally employs heuristic strategies to overcome noise and remain near the optimal path [piro2022optimal]. A detailed discussion of such strategies lies beyond the scope of this work.
[…]
@@ IV.5 Example 4: Snapshot of 2D Turbulence
[…]
sec:OPC_D
Next, we apply the numerical scheme described above to solve the optimal control equations for the steady background flow corresponding to a snapshot of two-dimensional turbulence, see Section LABEL:sec:Ex_D. The start and end points are chosen as $\bm{r}_{A}=(1.2,5.5)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Since the optimal trajectory only satisfies the local extremum condition, the resulting equations may, in general, admit multiple solutions. To illustrate this, Fig. LABEL:fig:6 presents two distinct optimal trajectories (shown as white lines) corresponding to different initial heading angles $\theta_{0}$ and total travel times $T$:
@@
[…]
Figure LABEL:fig:6 also illustrates the sensitivity of the faster trajectory to small perturbations in the initial control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minute variations in either of these quantities lead to trajectories that deviate significantly from the optimal path and ultimately fail to reach the target. This observation is further supported by the analysis of FTLEs presented in Fig. LABEL:fig:7, where the maximal FTLE attains positive values, confirming that the corresponding trajectory is unstable on the considered time scales. The observed sensitivity to small perturbations reflects an instability in the optimal control dynamics, associated with the agent’s self-propulsion activity [biferale2019zermelo].
@@
[…]
@@ IV.6 Example 5: 2D Turbulence
[…]
sec:OPC_E
In our final example, we consider a background velocity field corresponding to a fully time-dependent two-dimensional turbulent flow, see Section LABEL:sec:Ex_D. The starting and ending points are set to $\bm{r}_{A}=(1.4,3.0)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Figure LABEL:fig:8 presents the optimal trajectory (depicted as a white line), which corresponds to the initial heading angle $\theta_{0}$ and total travel time $T$ given by
@@
[…]
This trajectory represents one of potentially many solutions to the optimal control equations, reflecting the existence of multiple local extrema in the optimization landscape. While we cannot prove that the presented solution represents the absolute minimum travel time (i.e., the global optimum), it has been selected for illustration since it yields the shortest travel time among all optimal trajectories identified in our study. As in the previous example, Fig. LABEL:fig:8 highlights the sensitivity of the optimal trajectory to small perturbations in the initial conditions – both in the control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minor deviations in these parameters lead to trajectories that fail to reach the target, emphasizing the instability and complex structure of the underlying dynamical system. The positive value of the maximal FTLE, shown in Fig. LABEL:fig:9, further confirms the trajectory’s instability over the characteristic travel-time scale.
@@ IV.7 Summary
The optimal control framework provides a systematic method for determining time-minimizing trajectories between two points in a fluid flow for an active agent. In simple background flows, these trajectories are robust to small perturbations in initial position and steering direction, making them suitable for practical navigation. Once the initial steering angle $\theta_{0}$ is set, the agent can autonomously adjust its heading according to Eq. (LABEL:eq:angle), provided it can measure the local velocity field gradients. In more complex flows, finding optimal trajectories becomes considerably more difficult. The resulting paths often exhibit sensitivity to initial conditions and become unstable, limiting the direct application of a naive optimal control strategy in realistic environments. Nevertheless, such trajectories remain valuable. First, the optimal paths reveal how an agent can exploit the structure of the ambient flow to minimize travel time, thereby providing fundamental insight into efficient locomotion strategies. When supplemented with heuristic mechanisms designed to mitigate the influence of environmental noise and modeling uncertainties, these theoretically optimal solutions can still guide the construction of robust navigation algorithms [piro2022optimal]. Second, the optimal trajectories serve as valuable benchmarks: they establish theoretical lower bounds on travel time and offer reference solutions for evaluating the performance of alternative, stability-oriented navigation methods.
[…]
@@ V.1 Action-Value methods: Q-learning
The agent collects rewards along its way, and the goal is to find the optimal policy $\pi^{*}(a|s)$ that maximizes the expected cumulative reward over an episode. Action-value methods focus on learning the optimal function $q_{*}(s,a)$, which satisfies the Bellman equation:
[…]
@@
$label{eq:Q-update}\begin{aligned} Q(s_{t},a_{t})&\leftarrow Q(s_{t},a_{t})+\alpha\Big[\,r_{t+1}\\ &\quad+\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\,\Big],\end{aligned}$ (34)
@@
If the transition from the equation (LABEL:eq:Bellman) to the update rule (LABEL:eq:Q-update) appears unclear, Appendix LABEL:sec:appA provides a helpful explanation of the general idea.
@@ V.2 Policy-Gradient methods: Actor-Critic
where $\alpha_{p}$ denotes the learning rate. The core difficulty lies in computing the gradient $\nabla_{\bm{p}}J(\bm{p})$, since it depends not only on the actions selected by the policy but also on the distribution of states visited during interaction with the environment.
Fortunately, this gradient can be expressed analytically thanks to the policy gradient theorem, which provides a tractable form for optimization, see Appendix LABEL:sec:appB. According to the theorem,
[…]
@@
A notable improvement to REINFORCE involves introducing a baseline term $b(s)$, which helps reduce the variance of the gradient estimate without introducing bias. This modification is justified by the observation that
@@
since $\sum_{a}\pi(a|s,\bm{p})=1$. Hence, one can replace $q_{\pi}(s,a)$ with $q_{\pi}(s,a)-b(s)$ in the policy gradient theorem (LABEL:eq:theorem), leading to the modified update rule
@@ 5. $s\leftarrow s^{\prime}$
[…]
@@ V.3 States, Actions, Rewards
The state space is constructed by discretizing the spatial domain $L\times L$ into square tiles or cells $s_{i}$, with $i=1,2,\dots,N_{s}$, each of size $\delta\times\delta$. We set $\delta=L/N_{tile}$, so that in total we have $N_{s}=N_{tile}^{2}$ states. Roughly saying, this coarse graining implies that the agent resolves its position only up to the cell size $\delta$. The agent is allowed to change its steering direction only at discrete time intervals of duration $\Delta t=\delta/(2u_{0})$, where $u_{0}$ is taken to be of the order of the maximum flow velocity. The action space is also discretized: the control variable $\theta_{t}$ can take one of $N_{a}=8$ equally spaced angles $\theta_{j}=-\pi+2\pi j/N_{a}$ with $j=1,2,\dots,N_{a}$. During each time step of duration $\Delta t$, the agent’s trajectory is advanced via numerical integration of the governing equation of motion (LABEL:eq:motion) with the chosen steering angle held fixed.
@@
[…]
@@
where $p_{ij}$ is the $(i,j)$-th element of an $N_{s}\times N_{a}$ parameter matrix. Thus, each state-action pair $(s_{i},a_{j})$ is associated with a scalar parameter $p_{ij}$ that determines the probability of selecting action $a_{j}$ when the agent is in state $s_{i}$. For the state-value function, we also choose one of the most common parameterizations,
@@
where $w_{i}$ is the $i$-th component of a parameter vector $\bm{w}$ of length $N_{s}$. This representation assigns one free parameter to each state, keeping the function approximation simple. With these choices, the update rules (V.2) and (V.2) for a one-step actor-critic method take the form:

### v2

@@ IV.4 Example 3: Taylor-Green vortices
Next, we consider the Taylor-Green flow, see Section III.3. Now, the optimal control equations must be solved numerically. To define the problem over a fixed integration interval, it is convenient to introduce a rescaled time variable $\tau=t/T$, such that $\tau\in[0,1]$. Using this transformation, the governing equations take the form
$d\hat{X}_{\tau}/d\tau=u_{0}T\sin(k\hat{X}_{\tau})\cos(k\hat{Y}_{\tau})+V_{0}T\cos(\hat{\theta}_{\tau}),$ (21)
$d\hat{Y}_{\tau}/d\tau=-u_{0}T\cos(k\hat{X}_{\tau})\sin(k\hat{Y}_{\tau})+V_{0}T\sin(\hat{\theta}_{\tau}),$ (22)
$d\hat{\theta}_{\tau}/d\tau=u_{0}kT\big[\cos(k\hat{X}_{\tau})\cos(k\hat{Y}_{\tau})\sin(2\hat{\theta}_{\tau})$ (23)
$+\sin(k\hat{X}_{\tau})\sin(k\hat{Y}_{\tau})\big],$
where $\hat{\bm{R}}_{\tau}\equiv\bm{R}_{t}$, $\hat{\theta}_{\tau}\equiv\theta_{t}$ and the boundary conditions are $(\hat{X}_{0},\hat{Y}_{0})=(x_{A},y_{A})$ and $(\hat{X}_{1},\hat{Y}_{1})=(x_{B},y_{B})$.
@@
The system of equations can be numerically integrated starting from the initial position $\bm{r}_{A}$ for a prescribed initial steering angle $\theta_{0}$ and total travel time $T$. However, these two parameters are unknown and must be adjusted so that the final position $(\hat{X}_{1},\hat{Y}_{1})$ matches the desired target point $\bm{r}_{B}$. To achieve this, we employ a shooting method followed by fine-tuning optimization. Initially, approximate values of $\theta_{0}$ and $T$ are chosen to bring the trajectory close to the target. These parameters are then refined to minimize the discrepancy
${\mathcal{L}}(\theta_{0},T)=(\hat{X}_{1}-x_{B})^{2}+(\hat{Y}_{1}-y_{B})^{2}$ (24)
@@
between the simulated and desired endpoints. For optimization, we use gradient- and Hessian-based methods. In particular, our implementation first applies the Adam optimizer [kingma2014adam], followed by the BFGS algorithm [nocedal2006numerical]. Efficient computation of the loss gradients requires the forward integration scheme to support automatic differentiation [baydin2018automatic]. Moreover, the integration must be performed with high numerical accuracy. Convergence tests with progressively reduced tolerances are essential to ensure that the solution is independent of the integration precision.
Figure 4: Optimal control solution (white line) for the Taylor-Green flow, see Sections III.3. Perturbed trajectories are shown in cyan and blue for the initial control angle $\theta_{0}\pm\pi/9000$, and in magenta and red for the initial position error $\Delta=10^{-3}$ in $x_{A}$ and $y_{A}$.
For the numerical parameters specified in Section III.3, we obtain:
$\theta_{0}\approx 1.56079,\quad T\approx 12.2038.$ (25)
@@
The corresponding optimal trajectory is shown in Fig. 4 by a white line. The figure also illustrate the system’s sensitivity by depicting trajectories perturbed by small deviations in the initial heading $\theta_{0}\pm\pi/9000$ and by displacements $\Delta=10^{-3}$ in the initial $x_{A}$ and $y_{A}$ positions. Although these perturbations are considerably smaller than those examined in previous examples, the resulting trajectories fail to reach the target $\bm{r}_{B}$.
@@
To make a more quantitative stability analysis of the optimal trajectory, we come back to the original time variable $t$ and consider the evolution of infinitesimal perturbations,
$(\delta X_{t},\delta Y_{t},\delta\theta_{t})^{T}=\mathbb{W}(t)(\delta X_{0},\delta Y_{0},\delta\theta_{0})^{T},$ (26)
@@
where $\mathbb{W}(t)$ is the linearized evolution matrix that can be obtained numerically. A polar decomposition $\mathbb{W}(t)=\mathbb{V}(t)\mathbb{R}(t)$ allows us to separate the effects of rotation, represented by the orthogonal matrix $\mathbb{R}(t)$, from those of stretching, characterized by the positive-definite matrix $\mathbb{V}(t)$. The eigenvalues $h_{i}(t)$ of $\mathbb{V}(t)$ (equal to singular values of $\mathbb{W}(t)$) quantify the magnitude of stretching along different directions in phase space and are used to compute the finite-time Lyapunov exponents (FTLEs), defined as
@@
The corresponding results are presented in Fig. 5. The presence of a positive maximal FTLE indicates that the optimal trajectory is unstable with respect to small perturbations on the time scales needed to reach the target. This instability limits the practical applicability of the pre-computed optimal control strategy, as minor deviations in initial conditions lead to significant trajectory divergence. Moreover, the resulting sensitivity makes it difficult to find suitable initial guess for parameters $\theta_{0}$ and $T$, that lead to a successful solution of the optimization problem and finding optimal trajectories that reach the target. Despite these challenges, computing optimal trajectories can still be practically valuable if the agent additionally employs heuristic strategies to remain near the optimal path [piro2022optimal]. A detailed discussion of such strategies lies beyond the scope of this work.
Figure 5: Normalized FTLEs $\lambda_{i}T$ for the optimal trajectory found in the Taylor-Green flow.
@@ IV.5 Example 4: Snapshot of 2D Turbulence
Next, we apply the numerical scheme described above to solve the optimal control equations for the steady background flow corresponding to a snapshot of two-dimensional turbulence, see Section III.4. The start and end points are chosen as $\bm{r}_{A}=(1.2,5.5)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Since the optimal trajectory only satisfies the local extremum condition, the resulting equations may, in general, admit multiple solutions. To illustrate this, Fig. 6 presents two distinct optimal trajectories (shown as white lines) corresponding to different initial heading angles $\theta_{0}$ and total travel times $T$:
@@
Figure 6: Two optimal control solutions (white lines) for the snapshot of 2d turbulence, see Sections III.4. Perturbed trajectories for the lower path are shown in cyan and blue for the initial control angle $\theta_{0}\pm\pi/9000$, and in magenta and red for the initial position error $\Delta=10^{-3}$ in $x_{A}$ and $y_{A}$.
Figure 6 also illustrates the sensitivity of the faster trajectory to small perturbations in the initial control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minute variations in either of these quantities lead to trajectories that deviate significantly from the optimal path and ultimately fail to reach the target. This observation is further supported by the analysis of FTLEs presented in Fig. 7, where the maximal FTLE attains positive values, confirming that the corresponding trajectory is unstable on the considered time scales. The observed sensitivity to small perturbations reflects an instability in the optimal control dynamics, associated with the agent’s self-propulsion activity [biferale2019zermelo].
@@
Figure 7: Normalized FTLEs $\lambda_{i}T$ for the optimal trajectory found for the snapshot of 2d turbulence.
@@ IV.6 Example 5: 2D Turbulence
In our final example, we consider a background velocity field corresponding to a fully time-dependent two-dimensional turbulent flow, see Section III.4. The starting and ending points are set to $\bm{r}_{A}=(1.4,3.0)$ and $\bm{r}_{B}=(5.2,1.8)$, respectively. Figure 8 presents the optimal trajectory (depicted as a white line), which corresponds to the initial heading angle $\theta_{0}$ and total travel time $T$ given by
@@
Figure 8: Optimal control solution (white line) for the full time-dependent 2d turbulent flow, see Sections III.4. Perturbed trajectories are shown in cyan and blue for the initial control angle $\theta_{0}\pm\pi/9000$, and in magenta and red for the initial position error $\Delta=10^{-3}$ in $x_{A}$ and $y_{A}$.
This trajectory represents one of potentially many solutions to the optimal control equations, reflecting the existence of multiple local extrema in the optimization landscape. While we cannot prove that the presented solution represents the absolute minimum travel time (i.e., the global optimum), it has been selected for illustration since it yields the shortest travel time among all optimal trajectories identified in our study. As in the previous example, Fig. 8 highlights the sensitivity of the optimal trajectory to small perturbations in the initial conditions – both in the control parameter $\theta_{0}$ and the starting position $\bm{r}_{A}$. Even minor deviations in these parameters lead to trajectories that fail to reach the target, emphasizing the instability and complex structure of the underlying dynamical system. The positive value of the maximal FTLE, shown in Fig. 9, further confirms the trajectory’s instability over the characteristic travel-time scale.
@@ IV.7 Summary
The optimal control framework provides a systematic method for determining time-minimizing trajectories between two points in a fluid flow for an active agent. In simple background flows, these trajectories are robust to small perturbations in initial position and steering direction, making them suitable for practical navigation. Once the initial steering angle $\theta_{0}$ is set, the agent can autonomously adjust its heading according to Eq. (14), provided it can measure the local velocity field gradients. In more complex flows, finding optimal trajectories becomes considerably more difficult. The resulting paths often exhibit sensitivity to initial conditions and become unstable, limiting the direct application of a naive optimal control strategy in realistic environments. Nevertheless, such trajectories remain valuable. First, the optimal paths reveal how an agent can exploit the structure of the ambient flow to minimize travel time, thereby providing fundamental insight into efficient locomotion strategies. When supplemented with heuristic mechanisms designed to mitigate the influence of environmental noise and modeling uncertainties, these theoretically optimal solutions can still guide the construction of robust navigation algorithms [piro2022optimal]. Second, the optimal trajectories serve as valuable benchmarks: they establish theoretical lower bounds on travel time and offer reference solutions for evaluating the performance of alternative, stability-oriented navigation methods.
Figure 9: Normalized FTLEs $\lambda_{i}T$ for the optimal trajectory found for the full time-dependent 2d turbulent flow.
@@ V.1 Action-Value methods: Q-learning
The agent collects rewards along its way, and the goal is to find the optimal policy $\pi^{*}(a|s)$ that maximizes the expected cumulative reward over an episode. Action-value methods focus on learning the corresponding function $q_{*}(s,a)$, which satisfies the Bellman equation:
$q_{*}(s,a)=\mathbb{E}\!\left[r_{t+1}+\max_{a^{\prime}}q_{*}(s_{t+1},a^{\prime})|s_{t}=s,a_{t}=a\right].$ (33)
@@
$Q(s_{t},a_{t})$ $\leftarrow Q(s_{t},a_{t})+\alpha\Big[\,r_{t+1}$ (34)
$\quad+\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\,\Big],$
@@
If the transition from the equation (33) to the update rule (34) appears unclear, Appendix A provides a helpful explanation of the general idea.
@@ V.2 Policy-Gradient methods: Actor-Critic
where $\alpha_{p}$ denotes the learning rate. The core difficulty lies in computing the gradient $\nabla_{\bm{p}}J(\bm{p})$.
Fortunately, this gradient can be expressed analytically thanks to the policy gradient theorem, which provides a tractable form for optimization, see Appendix B. According to the theorem,
$\nabla_{\bm{p}}J(\bm{p})\propto\sum_{s}\mu(s)\sum_{a}q_{\pi}(s,a)\nabla_{\bm{p}}\pi(a|s,\bm{p}),$ (37)
@@
A notable improvement to the reinforce algorithm involves introducing a baseline term $b(s)$, which helps reduce the variance of the gradient estimate without introducing bias. This modification is justified by the observation that
@@
since $\sum_{a}\pi(a|s,\bm{p})=1$. Hence, one can replace $q_{\pi}(s,a)$ with $q_{\pi}(s,a)-b(s)$ in the policy gradient theorem (37), leading to the modified update rule
@@ 5. $s\leftarrow s^{\prime}$
Figure 10: Example of steady vortex flow, see Section III.1. (a) Total reward, averaged over a sliding window of $250$ episodes, plotted against the number of training episodes. (b) Trajectories generated by the Q-learning (red) and one-step actor-critic (blue) agents, shown alongside the optimal control trajectory (white). (c) Arrival-time histograms for the trained agents.
@@ V.3 States, Actions, Rewards
The state space is constructed by discretizing the spatial domain $L\times L$ into square tiles or cells $s_{i}$, with $i=1,2,\dots,N_{s}$, each of size $\delta\times\delta$. We set $\delta=L/N_{tile}$, so that in total we have $N_{s}=N_{tile}^{2}$ states. Roughly saying, this coarse graining implies that the agent resolves its position only up to the cell size $\delta$. The agent is allowed to change its steering direction only at discrete time intervals. The duration between them is $\Delta t=\delta/(2u_{0})$, where $u_{0}$ is taken to be of the order of the maximum flow velocity. The action space is also discretized: the control variable $\theta_{t}$ can take one of $N_{a}=8$ equally spaced angles $\theta_{j}=-\pi+2\pi j/N_{a}$ with $j=1,2,\dots,N_{a}$. During each time step of duration $\Delta t$, the agent’s trajectory is advanced via numerical integration of the governing equation of motion (1) with the chosen steering angle held fixed.
@@
Figure 11: Example of steady sink and time-varying vortex flow, see Section III.2. (a) Total reward, averaged over a sliding window of $250$ episodes, plotted against the number of training episodes. (b) Trajectories generated by the Q-learning (red) and one-step actor-critic (blue) agents, shown alongside the optimal control trajectory (white). (c) Arrival-time histograms for the trained agents.
@@
where $p_{ij}$ is the $(i,j)$-th element of an $N_{s}\times N_{a}$ parameter matrix. Thus, each state-action pair $(s_{i},a_{j})$ is associated with a scalar parameter $p_{ij}$ that determines the probability of selecting action $a_{j}$ when the agent is in state $s_{i}$. Finally, for the state-value function we also choose one of the most common parameterizations,
@@
where $w_{i}$ is the $i$-th component of a parameter vector $\bm{w}$ of length $N_{s}$. This representation assigns one free parameter to each state, keeping the function approximation simple. With these choices, the update rules (V.2) and (V.2) for the one-step actor-critic method take the form: